OpenStudy (anonymous):

The polygon Angle-Sum Theorem states:The sum of the measures of the angles of an n-gon is a.)n-2/180 b.)(n-1)180 c.)180/n-1 d.)(n-2)180

OpenStudy (solomonzelman):

we have used this theorem many times, don't you recall it? Which one do you think it is?

OpenStudy (anonymous):

would it be where (n-2)180??

OpenStudy (solomonzelman):

Yes,

OpenStudy (anonymous):

so id would be d??

OpenStudy (anonymous):

so it would be d??

OpenStudy (solomonzelman):

Yes, it is D.

OpenStudy (anonymous):

okay can you help me on my last question??

OpenStudy (solomonzelman):

you are going to have a good Sunday... you are done already :D

OpenStudy (solomonzelman):

sure, Y-not

OpenStudy (anonymous):

Find the missing values of the variables. The diagram is not drawn to scale a.)x=124;y=125 b.)x=56;y=114 c.)x=114;y=56 d.)x=56;y=124

OpenStudy (anonymous):



OpenStudy (solomonzelman):

124 is that angle next to y?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

1) 124 and y are supplementary angles, saying that 124+y=180 Can you solve for ? (this is just step 1)

OpenStudy (solomonzelman):

I meant to say, can you solve for y (knowing that 180=y+124)

OpenStudy (anonymous):

so you take minus 124 from both sides so you you'd get y=56??

OpenStudy (anonymous):

oh and then x would be 114??

OpenStudy (solomonzelman):

yes, correct. y=56 And then we can apply our favorite rule that your lesson is (probably) about. We have a 4-gon

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

\(\large\color{blue}{ (4-2)180=125+65+56+x }\) \(\large\color{blue}{360=246+x }\) \(\large\color{blue}{x=114 }\)

OpenStudy (solomonzelman):

You are totally correct, I was just checking....

OpenStudy (anonymous):

Thank you so much for your help :)

OpenStudy (solomonzelman):

Have fun :)